The radioactive isotope 198Au has a half-life of 64.8h. A sample containing this isotope has an initial activity of 40.0 Bq. Calculate the number of nuclei that decay in the time interval between in 2h.
1
Expert's answer
2021-02-18T18:38:09-0500

"t_{1/2}=64.8 \;h = 64.8 \times 3600 = 233280 \;sec"

Initial activity = 40.0 Bq

"t = 2 \;h \\

λ = \frac{ln(2)}{t_{1/2}} \\

= \frac{ln(2)}{233280} \\

= 2.97 \times 10^{-6} \;s^{-1} \\

A_0 = λN_0 \\

N_0 = \frac{A_0}{λ} \\

= \frac{40}{2.97 \times 10^{-6}} \\

= 1.346 \times 10^7 \\

N = N_0e^{-λt}"

"= 1.346 \times 10^7 \times e^{-2.97 \times 10^{-6} \times 2 \times 3600} \\

= 1.346 \times 10^7 \times e^{-0.02138} \\

= 1.346 \times 10^7 \times 0.9788 \\

= 1.317 \times 10^7"

The number of nuclei that decay

"= N_0 – N \\

= 1.346 \times 10^7 - 1.317 \times 10^7 \\

= 0.029"